OpenStudy (anonymous):

In a scene in an action movie, a stuntman jumps from one building to the top of another building 4.0 m away. After a running start he leaps at a velocity of 5.0 m/s at an angle of 15 degrees with respect to the flat roof. Will he make it to the other roof, which is 2.5 m shorter than the building he jumps from?

OpenStudy (anonymous):

what level is this question for? e.g. high school, uni?

OpenStudy (anonymous):

\[[5\sin(15)]t-(1/2)[10\times t ^{2}]=-2.5\] i have taken the value of g as \[g= 10 m/s ^{2}\]the value of t i get in this equation, i would substitute in the following \[5\cos(15)t=x\] if i get the value of x as \[x \ge 4m\]then he will make it otherwise he won't.